OpenStudy (anonymous):

what is the slope of a line x=5

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

undefined

OpenStudy (anonymous):

slope of any line is y=mx+c and so here in the given equation there no y term tthhis means mhas to tend to infinity ...which we consider as undefined.

OpenStudy (anonymous):

slope is infinity because the equation can be re-written as 0y +x - 5 = 0 0y = -x + 5 (y = mx +c form) slope = -1 / 0 which is infinity

OpenStudy (anonymous):

infinity...